# A general comparison principle for the pluripotential complex Monge-Ampère flow

### v1

@@ 1. Introduction
Let $dV$ be a Lebesgue measure on $\mathbb{C}^{n}$. Guedj, Lu and Zeriahi [GLZ21-2] proved the existence and uniqueness of the solution to the Cauchy-Dirichlet problem with a measure $\mu$ and a boundary data $h$ satisfying
@@
Finally, let us give the historical background of the complex Monge-Ampère flows before we finish the introduction. In 1985, Cao [Cao85] reproved the Yau’s solution to the Calabi conjecture and existence of Kähler-Einstein metric when the first Chern class is negative or zero by using the Kähler-Ricci flow. Since then, the Kähler-Ricci flow has been an important topic in itself with various research subjects, including its behavior on Fano manifolds [PSSW, PS06, Sz10], extension on non-Kähler settings [Gi11, TW15], and so on.
@@
For the complex Monge-Ampère flows, studying the weak solutions of degenerate equations has been a main interest for its geometrical application, especially for the analytic minimal model program proposed by Song and Tian [ST17]. Eyssidieux, Guedj and Zeriahi [EGZ15, EGZ16] first studied the viscosity approach. Later, Guedj, Lu and Zeriahi [GLZ20, GLZ21-2] introduced the pluripotential approach. They also proved in [GLZ21-1] that both approaches are equivalent under some assumptions. Recently, Dang [Dan24] also extended the pluripotential approach to the Chern-Ricci flow. We refer the reader to [To21, Da22] for the application of viscosity and pluripotential approaches. Besides its geometrical applications, degenerate complex Monge-Ampère flow has been also an interesting topic itself. We refer the reader to [DLT20, DP22] and [Do16, Do17b] for various generalizations of viscosity approach and more general initial data, respectively.
Organization. In Section $2$, we recall the definition of parabolic Monge-Ampère operator in [GLZ21-2]. We also prove several useful lemmas about the operator. In Section $3$, we first prove some lemmas about mixed type inequalities and property of supersolutions. Then we give the proof of Theorem 1.2 in Theorem 3.8 by using these lemmas. With this result, we also prove the uniqueness of the weak solution to the Cauchy-Dirichlet problem and study the long-term behavior of the solution.
Acknowledgements. I would like to thank my advisor, Ngoc Cuong Nguyen, for providing invaluable guidance and support. I would like to thank Quan Tuan Dang for helpful discussion on the proof of the Theorem 3.8 and to Hyunsoo Ahn for a careful reading of the earlier version of its proof. I also thank to Do Hoang Son for useful comments.
@@ 2. Parabolic Monge-Ampère Operator
For sufficiently small $\varepsilon_{2}>0$, $u_{\varepsilon_{2}}$ is defined on $(0,T)\times\Omega^{\prime}$ for some open set $\Omega^{\prime}$ satisfying $K\Subset\Omega^{\prime}\Subset\Omega$. Moreover, one can check that $u_{\varepsilon_{2}}\in\mathcal{P}((0,T)\times\Omega^{\prime})\cap C((0,T)\times\Omega^{\prime})$ and $u_{\varepsilon_{2}}\downarrow u$ as $\varepsilon_{2}\rightarrow 0$. Since $v$ is (jointly) upper semi-continuous by [GLZ21-2, Proposition 1.4], $\{v<u_{\varepsilon_{2}}-\varepsilon_{1}\}\cap((0,T^{\prime})\times\Omega^{\prime})$ is an open subset of $\Omega_{T}$ which increases to $\{v<u-\varepsilon_{1}\}\cap((0,T^{\prime})\times\Omega)$ as $\varepsilon_{2}\rightarrow 0$. By taking $\varepsilon_{1}\rightarrow 0$ and $T^{\prime}\rightarrow T$, one can see that $\{v<u\}$ is a Borel subset of $\Omega_{T}$.
@@
It follows from Lemma 2.2 that each side of (2.7) is a Borel measurable function on $(0,T)$. By integrating each side with respect to $dt$ for $t\in(0,T)$,
@@ 3.3. Main result
The idea is to prove that $t_{1}>0$ as in the left figure, and then compute the size of the set $E\cap((t_{1}-\delta,t_{1}+\delta)\times\Omega)$ which is described in the right figure. We will show that it is actually an empty set, which contradicts to the definition of $t_{1}$.
@@
for all $t,s\in[t_{1}-\tilde{\varepsilon},t_{1}+\tilde{\varepsilon}$] and $z\in\Omega$. It follows from the assumption (c) that there exists $C_{2}>0$ such that $u(t,z)-C_{2}t^{2}$ and $v(t,z)-C_{2}t^{2}$ is concave in $t_{1}-\tilde{\varepsilon},t_{1}+\tilde{\varepsilon}$] for all $z\in\Omega$. Let us define $C_{3}:=\max\{C_{1},C_{2}\}$, so that we have
@@
for some $B>0$ in $\Omega_{S}$. We will remove this assumption at the end of the proof. For $s>0$ near $1$ we set, for $(t,z)\in\Omega_{S}$,
@@
and apply Lemma 3.6. Note that (i) follows from Lemma 3.3. To prove (ii), it suffices to show that for each open $U\Subset\Omega$, there exists $A_{0}>0$ such that

### v2

@@ 1. Introduction
Let dVdV be a Lebesgue measure on $\mathbb{C}^{n}$. Guedj, Lu and Zeriahi [GLZ21-2] proved the existence and uniqueness of the solution to the Cauchy-Dirichlet problem with a measure $\mu$ and a boundary data $h$ satisfying
@@
Finally, let us give the historical background of the complex Monge-Ampère flows before we finish the introduction. In 1985, Cao [Cao85] reproved the Yau’s solution to the Calabi conjecture and existence of Kähler-Einstein metric when the first Chern class is negative or zero by using the Kähler-Ricci flow. Since then, the Kähler-Ricci flow has been an important topic in itself with various research subjects, and we refer the including its behavior on Fano manifolds [PSSW, PS06, Sz10], extension on non-Kähler settings [Gi11, TW15], and so on. We refer the reader to survey papers [SW13, To18] for more details.
@@
For the complex Monge-Ampère flows, studying the weak solutions of degenerate equations has been a main interest for its geometrical application, especially for the analytic minimal model program proposed by Song and Tian [ST17]. Eyssidieux, Guedj and Zeriahi [EGZ15, EGZ16] first studied the viscosity approach. Later, Guedj, Lu and Zeriahi [GLZ20, GLZ21-2] introduced the pluripotential approach. They also proved in [GLZ21-1] that both approaches are equivalent under some assumptions. Recently, Dang [Dan24] also extended the pluripotential approach to the Chern-Ricci flow to study the question proposed by Tosatti and Weinkove [TW22]. We refer the reader to [To21, Da22] for the application of viscosity and pluripotential approaches. Besides its geometrical applications, degenerate complex Monge-Ampère flow has been also an interesting topic itself. We refer the reader to [DLT20, DP22] and [Do16, Do17b] for various generalizations of viscosity approach and more general initial data, respectively.
Organization. In Section 22, we recall the definition of parabolic Monge-Ampère operator in [GLZ21-2]. We also prove several useful lemmas about the operator. In Section 33, we first prove some lemmas about mixed type inequalities and property of supersolutions. Then we give the proof of Theorem 1.2 in Theorem 3.8 by using these lemmas. With this result, we also prove the uniqueness of the weak solution to the Cauchy-Dirichlet problem and study the long-term behavior of the solution.
Acknowledgements. I would like to thank my advisor, Ngoc Cuong Nguyen, for providing invaluable guidance and support. I would like to thank Quan Tuan Dang for helpful discussion on the proof of the Theorem 3.8 and to Hyunsoo Ahn for a careful reading of the earlier version of its proof. I also thank to Do Hoang Son and Sławmoir Dinew for helpful comments.
@@ 2. Parabolic Monge-Ampère Operator
For sufficiently small $\varepsilon_{2}>0$, $u_{\varepsilon_{2}}$ is defined on $(0,T)\times\Omega^{\prime}$ for some open set $\Omega^{\prime}$ satisfying $K\Subset\Omega^{\prime}\Subset\Omega$. Moreover, one can check that $u_{\varepsilon_{2}}\in\mathcal{P}((0,T)\times\Omega^{\prime})\cap C((0,T)\times\Omega^{\prime})$ and $u_{\varepsilon_{2}}\downarrow u$ as $\varepsilon_{2}\rightarrow 0$. Since $v$ is (jointly) upper semi-continuous by [GLZ21-2, Proposition 1.4], $\{v<u_{\varepsilon_{2}}-\varepsilon_{1}\}\cap((0,T^{\prime})\times\Omega^{\prime})$ is an open subset of $\Omega_{T}$ which decreases to $\{v<u-\varepsilon_{1}\}\cap((0,T^{\prime})\times\Omega)$ as $\varepsilon_{2}\rightarrow 0$. By taking $\varepsilon_{1}\rightarrow 0$ and $T^{\prime}\rightarrow T$, one can see that $\{v<u\}$ is a Borel subset of $\Omega_{T}$.
@@
It follows from Lemma 2.2 that each side of (2.7) is a Borel measurable function on $(0,T)$. By integrating each side with respect to dtdt for $t\in(0,T)$,
@@ 3.2. Boundary behavior of supersolutions
The idea is to prove that $t_{1}>0$ as in the left figure, and then compute the size of the set $E\cap((t_{1}-\delta,t_{1}+\delta)\times\Omega)$ which is described in the right figure, where $\delta>0$ will be chosen later. We will show that it is actually an empty set, which contradicts to the definition of $t_{1}$.
@@
for all $t,s\in[t_{1}-\tilde{\varepsilon},t_{1}+\tilde{\varepsilon}$] and $z\in\Omega$. It follows from the assumption (c) that there exists $C_{2}>0$ such that $u(t,z)-C_{2}t^{2}$ and $v(t,z)-C_{2}t^{2}$ are concave in $t_{1}-\tilde{\varepsilon},t_{1}+\tilde{\varepsilon}$] for all $z\in\Omega$. Let us define $C_{3}:=\max\{C_{1},C_{2}\}$, so that we have
@@
for some $B>0$ in $\Omega_{S}$. We will remove this assumption at the end of the proof. For $s>0$ near 11 we set, for $(t,z)\in\Omega_{S}$,
@@
and apply Lemma 3.6. Note that (i) follows from Lemma $3.$ To prove (ii), it suffices to show that for each open $U\Subset\Omega$, there exists $A_{0}>0$ such that